two vectors f1=250 N due east and f2= 250 N due north have common initial point. find f2-f1 and it’s direction?

Grade:11

1 Answers

Eshan
5 years ago
\vec{F_2}-\vec{F_1}=\vec{F_2}+(-\vec{F_1})
=250due north +250due west.

Hence the resultatn is in North west with a magnitude of\sqrt{250^2+250^2}=250\sqrt{2}